A 100 gram apple is released from rest 500 m above the ground. Due to the combined forces of air resistance and gravity, it has a speed of only 40 m/s when it reaches the ground. How much work must have been done on the apple by the force of air resistance?

(A) 80 J

(B) 420 J

(C) 500 J

(D) 580 J




1
Expert's answer
2021-03-25T19:29:02-0400

"\Delta W=\frac{mv^2}{2}-mgh=\frac{0.1\cdot 40^2}{2}-0.1\cdot10\cdot 500=-420\ (J)"


"(B)\ 420\ J" . Answer